# Inverse Constitutional AI: Compressing Preferences into Principles

### v1

@@ 4.1 Synthetic data as a proof of concept
This dataset is based on preferences according to orthogonal principles, intended to be on average neither supported nor opposed by humans or language models. In particular, we create a dataset based on three principles: “prefer cats over dogs”, “prefer green over blue color”, and “select lemon over raspberry ice-cream”. We generate 10 pairs per principle, overall 30 pairs.
Aligned.
This dataset uses preferences generally accepted by humans and (especially) language models. Our dataset follows three principles: “select truthful over factually incorrect answers”, “select helpful over useless answers”, “select polite over impolite answers”. We again generate 10 pairs per principle, overall 30 pairs.
Unaligned.
This dataset flips the aligned dataset’s annotations, creating a dataset a default LLM annotator mostly disagrees with.
Figure 3 shows our results on synthetic data. We compare default annotators (prompted to select the “best” model) to constitutional annotators (prompted with a generated constitution). LLM annotators with generated constitutions reconstruct original annotations better in the orthogonal and unaligned datasets, and keep high agreement in the aligned case. These results indicate the constitutions capture helpful information about the preferences. Qualitatively, the generated constitutions [^6] The best, median and worst constitutions are available in full in Appendix C for all relevant experiments. often closely correspond to the principles described above.
@@ 4.2 Human-annotated AlpacaEval data
Figure 4: Results on AlpacaEval data. GPT-4o is capable of generating and using constitutions that notably improve agreement with unaligned preferences. We sample 65 training and 65 test samples from AlpacaEval’s human-annotated dataset, considering both aligned (original) and unaligned (flipped annotations) versions. GPT-4o generates constitutions for each version on the training set, evaluated on the test set using GPT-3.5-Turbo and GPT-4o annotators. In the aligned case, GPT-3.5-Turbo slightly increases its average agreement using a constitution while GPT-4o decreases its average agreement. This may be because GPT-4o is more aligned with human annotations than our average compressed constitution, while GPT-3.5-Turbo is less aligned and benefits from the constitution. In the unaligned case, GPT-4o’s agreement improves notably, while GPT-3.5-Turbo’s performance does not exceed random choice, indicating its limited ability to follow unaligned principles. Reported values and error bars are the mean and standard deviation over 6 random seeds. Example constitutions were picked for illustrative purposes.
Next, we test our approach on human-annotated texts, using the AlpacaEval dataset [^7] https://huggingface.co/datasets/tatsu-lab/alpaca_eval, license: CC-BY-NC-4.0[Dubois et al., 2023]. The dataset, used for the AlpacaEval leaderboard, contains about 650 data points cross-annotated by four annotators, with well-tested baseline AI annotators and evaluation tooling.
@@
Experimental setup. We randomly select two mutually exclusive training and test subsets, each with 65 annotated pairs (overall $\mathord{\sim}20\%$ of the dataset). Similar to the synthetic experiments, we derive an aligned and unaligned dataset. In the aligned version the annotations are based on the majority vote of the original 4 cross-annotations (ties broken randomly), while the annotations are flipped in the unaligned scenario, so that most humans (and LLMs) would disagree with them.
Results. Figure 4 shows the results in aligned and unaligned scenarios. Constitutions are generated on the training subset and results reported on the (unseen) test subset. In general, we observe that the improvement with constitutions are limited in the aligned scenario but more noticeable in the unaligned scenario. There appears to be a capability gap between GPT-3.5-Turbo and GPT-4o in the unaligned case: whereas the former does not improve beyond random with a constitution, the latter is able to notably outperform the default annotator and performs well above random 50% agreement.
@@
Finally, given GPT-3.5-Turbo’s limitations in the unaligned case in Fig. 4, we explore more generally how well our constitutions can transfer between models. To this end, we take the best performing constitution of the unaligned case on the training set and test how well models from Anthropic’s Claude family are able to use these constitutions on the test set. Shown in Fig. 5, the results indicate that this constitution transfers well to the Claude models–––better than to GPT-3.5-Turbo.
Figure 5: Transferability of constitutions: results of transferring a GPT-4o generated constitution to other model family (Claude). We use the highest-performing unaligned constitution on the training set, from experiments shown in the unaligned plot in Figure 4. We test two additional models from the Claude model family, Claude-3-Opus and Claude-3-Haiku. Both are able to use GPT-4o’s generated constitution to reconstruct the test set annotations effectively, albeit to a lower standard than GPT-4o. The reported values and error bars are the mean and standard deviation over 4 random seeds per annotator, all with the same constitution.
@@ 4.3 Individual preferences: Chatbot Arena data
We test our approach’s ability to generate personal constitutions for individual users using the Chatbot Arena Conversations dataset [^8] https://huggingface.co/datasets/lmsys/chatbot_arena_conversations, license: CC-BY-NC-4.0 [Zheng et al., 2023], which consists of 33k human-annotated preferences filtered by the authors to avoid personally identifiable information (PII). Unlike AlpacaEval, the dataset includes user-generated prompts in addition to user-generated preferences, which may be considered more expressive of user preferences than annotations on pre-existing prompts.
Experimental setup. We select two users exhibiting different preferences from the Chatbot Arena dataset (with 9 and 12 preference annotations respectively) and generate a separate constitution with 3 principles for each. To be better able to detect the effect of user-specific principles, we adapt our generation and annotation prompts to be less general, i.e. to generate constitutions more specific to the individual users and to follow the specific principles more closely rather than the model’s priors.
Results. Figure 6 shows the results of our experiments: the personal constitutions of each user work best for the user they were created for. Note that how well a personal constitution transfers from one user to another depends on how similar their preferences are. Constitutions of similar users may transfer well, whilst the contrast will be most pronounced between users with opposing preferences. Thus, the results here illustrate the use of personal constitutions, but may vary for any given pair of users.
Figure 6: Case-study of generating personal constitutions for anonymous Chatbot Arena users. Personal constitutions have the potential to help make LLM applications more helpful and customized to individual users’ preferences—in an interpretable way. We generate constitutions based on a single user’s annotations and check the constitutions’ ability to help re-construct the annotations of the same user and another user. For the selected users, we observe that generated personal constitutions appear to work best for original user and do not transfer perfectly to another user. Note that this effect will vary for other users depending on how different users’ preferences are. Examples of generated constitutions were chosen for illustrative purposes. Reported values and error bars are the mean and standard deviation over 6 random seeds of the entire generation-annotation pipeline, run with GPT-4o.
@@ 5 Related work
Our work focuses on deriving interpretable principles from human feedback data and using AI annotators to evaluate those principles. We build on work related to learning from human feedback, biases in feedback, interpretable preference models, prompt generation, and AI annotators.
@@
Learning from human feedback. Fine-tuning LLMs with human feedback has significantly contributed to the success of modern LLMs [Ouyang et al., 2022, Stiennon et al., 2020]. Typically, feedback is collected through pairwise comparisons of model outputs, training a reward model for fine-tuning, e.g.  using reinforcement learning from human feedback (RLHF) [Ouyang et al., 2022, Stiennon et al., 2020, Kaufmann et al., 2024] or direct preference optimization (DPO) [Rafailov et al., 2023]. Interpreting preference data is challenging since it generally lacks annotations of the underlying reasons. Additionally, the reward model is often a black-box neural network, making it hard to interpret as well. Our work aims to generate interpretable principles explaining the feedback data.
@@
Biases in human feedback. Identifying biases in human feedback data is crucial, as unintended biases are common [Hosking et al., 2024, Wu and Aji, 2023, Bansal et al., 2024, Sharma et al., 2023, Xu et al., 2023]. For example, Hosking et al. [2024] note a preference for assertiveness over truthfulness, while Wu and Aji [2023] highlight a bias towards grammatical correctness over factual accuracy. Bansal et al. [2024] highlight that feedback methods influence biases, e.g., annotators focus more on accuracy in pairwise ranking compared to rating feedback. Additionally, Sharma et al. [2023] observe a bias towards sycophantic outputs, where responses align with the user’s beliefs rather than the truth. While these studies provide valuable insights, most methods for detecting biases rely on specialized feedback data collection, making them challenging to apply to pre-existing data. Our work generates interpretable principles from existing preference data, which can be inspected to detect biases and provide insights into the underlying preferences.
@@
Interpretable preference models. There has been a growing interest in creating interpretable preference models, aiding in understanding behavior of AI systems. Go et al. [2024] create a compositional preference model based on 13 fixed features, similar to our constitutional principles. While they do not generate the constitution from data, they do create a regression model to weight them, which would be a promising extension to our approach. Petridis et al. [2024] propose a feedback-based constitution generation method relying on interactive tools, whereas our approach can be applied to standard preference datasets.
Prompt generation. LLM outputs can be guided by generating specific prompts. This relates closely to our work, where we create principles to steer outputs. Adversarial prompt generation aims to produce outputs contrary to the model’s training data, similar to our unaligned setting. Manual adversarial prompt generation, or ‘jailbreaking’, allows users to bypass safety constraints imposed during fine-tuning. This process can also be automated [Zou et al., 2023], generating adversarial prompts to attack a wide range of models. Li et al. [2024a] propose virtual tokens to steer outputs towards specific viewpoints, using a dataset of question responses to define these personas, unlike our approach based on pairwise comparisons and interpretable constitutions. Rodriguez et al. [2024] explore the use of LLMs to discover and classify user intent, which may help adapt model prompts.
@@
AI annotators. Due to the cost and time required for human feedback collection, AI annotators, or LLM-as-a-judge, have been proposed as a scalable alternative. Constitutional AI [Bai et al., 2022] uses LLMs with a set of principles for feedback. Due to strong alignment with human preferences through fine-tuning, LLMs can generalize from very general principles, such as “do what’s best for humanity” [Kundu et al., 2023], or even give feedback well-aligned with human preferences without any constitution [Zheng et al., 2023]. Our experiments show similar trends, where default LLM annotators align well with dataset annotations, even without a constitution. AI annotators can exhibit biases, partially overlapping with human biases [Chen et al., 2024], and inconsistencies in their judgments [Stureborg et al., 2024]. Examples include position bias (preferring the first output) [Zheng et al., 2023], verbosity bias (preferring longer responses) [Zheng et al., 2023], and self-enhancement or familiarity bias (preferring outputs similar to their own) [Panickssery et al., 2024, Stureborg et al., 2024]. Their proposed mitigation measures include trying both orderings and tying if inconsistent, with further explorations in later work [Dubois et al., 2024]. AlpacaEval [Li et al., 2024b] offers a set of well-validated AI annotators. Our work both builds upon and extends this line of research, proposing a new method to automatically generate an inspectable and modifiable constitution and using an LLM-as-a-judge to guide an LLM towards this constitution.
@@
Rule-based preference learning. Our work involves rule learning [Fürnkranz et al., 2012], aiming to develop descriptive or predictive rules, a method previously applied to preference learning [de Sá et al., 2011]. A common technique for rule learning is to first generate a set of candidate rules and then measure each rule’s support in a dataset, i.e. the fraction of data points that satisfy the rule [Fürnkranz et al., 2012, de Sá et al., 2011]. Our algorithm follows this approach but, in contrast to more traditional rule learning, generates rules as natural language sentences. These rules, though more ambiguous and requiring AI annotators for interpretation, are expressive, interpretable, and easy for non-experts to edit.
6 Limitations and discussion

### v2

@@ 4.1 Proof-of-concept: Synthetic data
This dataset is based on principles intended to be neither supported nor opposed by humans or language models on average. In particular, we create a dataset based on three principles: “prefer cats over dogs”, “prefer green over blue color”, and “select lemon over raspberry ice-cream”.
Aligned and unaligned.
The aligned dataset uses preferences generally accepted by humans and (especially) language models. Our dataset follows three principles: “select truthful over factually incorrect answers”, “select helpful over useless answers”, “select polite over impolite answers”. The unaligned dataset flips these annotations, creating a dataset that a default LLM annotator mostly disagrees with.
Figure 3: Results on synthetic data. Our constitutional annotators can reconstruct a variety of preferences using limited data and without fine-tuning. We demonstrate our algorithm’s adaptability on three synthetic datasets: one orthogonal to the base LLM’s learned preferences, one aligned with those preferences and one unaligned with them. We generate constitutions for each and report agreement with the original data of a default LLM annotator and a constitutional annotator (prompted with a constitution). Our constitutions notably improve agreement in the orthogonal and unaligned cases and retain high agreement in the aligned case, albeit with more variance. Our method’s ability to detect biases is illustrated by the example constitution in the unaligned case. Plots show mean and standard deviation (6 seeds) using GPT-3.5-Turbo.
Results. In Figure 3, we compare default annotators (prompted to select the “best” output) to constitutional annotators (prompted with a generated constitution). We find that constitutional annotators reconstruct original annotations better in the orthogonal and unaligned datasets, and keep high agreement in the aligned case (which offers limited room for improvement). These results indicate that the constitutions capture helpful information about the preferences. Qualitatively, the generated constitutions (see Appendix H) often closely correspond to the principles described above.
@@ 4.2 Human-annotated AlpacaEval data
Figure 4: Results on AlpacaEval data. GPT-4o generates and uses interpretable constitutions that match the performance of the default annotator on aligned preferences and notably increase agreement with unaligned preferences. Tested on aligned (original) and unaligned (flipped) versions of AlpacaEval, with GPT-4o generating constitutions which are then used by constitutional annotators backed by GPT-4o and GPT-3.5-Turbo. Note we can only expect significant improvement in the unaligned case, as discussed in the main text. The aligned case does not leave room for improvement over the default annotator, but allows us to gain new insights into the preferences expressed in the dataset. In the unaligned case, GPT-4o’s agreement improves notably, while GPT-3.5-Turbo’s performance does not exceed random choice, indicating its limited ability to follow unaligned principles. Plots show mean and standard deviation (6 seeds).
We test our approach on human-annotated texts using the AlpacaEval dataset (Dubois et al., 2023). The dataset, used for the AlpacaEval leaderboard, features 648 data points cross-annotated by four annotators, with well-tested baseline AI annotators and evaluation tooling. The dataset captures general human preferences, likely very similar to the preferences language models used in this work have been fine-tuned on. As a consequence, the default annotator (without a constitution) agrees strongly with the annotations, leaving little room for improvement. The goal of this experiment, then, is not to exceed the default annotator’s annotation performance on this dataset but to answer the following research questions: (Q1) Can constitutional annotators match the default annotator’s performance on the aligned dataset, while providing the benefit of interpretable and editable constitutions? (Q2) Is ICAI able to extract and follow principles that are exactly opposite to the aligned dataset, despite the underlying model’s biases? Note that most practical applications will be somewhere between the two extremes of the aligned and the unaligned scenario, allowing ICAI to increase the annotator’s performance while offering insights into the learned principles, along with the ability to inspect and modify them as needed.
@@
Experimental setup. For each seed, we randomly select mutually exclusive training and test subsets with 65 annotated pairs each. Constitutions are generated on the training subset and results reported on the (unseen) test subset. We derive an aligned dataset based on majority vote (ties broken randomly) and an unaligned dataset using flipped annotations.
Results. Figure 4 shows that constitutional annotators approximately match default annotator performance in the aligned scenario while using an easily interpretable constitution, answering (Q1) affirmatively. [^2] We observe a very slight improvement in GPT-3.5-Turbo’s annotations and a similarly small reduction in GPT-4o’s agreement. This may be explained by the GPT-3.5-Turbo model being less attuned to the preferences in the dataset, which can be alleviated with a constitution. In the case of GPT-4o, however, the constitutional annotator likely focuses on the highly compressed constitution, even in cases where the default annotator would have a more nuanced (but less interpretable) understanding of the underlying preferences. The unaligned dataset shows GPT-4o succeeding at following opposing principles, improving beyond the default annotator, while GPT-3.5-Turbo fails to do so despite using the same GPT-4o-generated constitutions. This answers (Q2) affirmatively for GPT-4o, revealing a capability gap between models. Since we evaluate the constitutions on an unseen test set, these results also demonstrate ICAI’s potential for annotation scaling, extracting a constitution from a small training set (65 preferences here) and applying it to new data.
@@
Constitution transferability. Given GPT-3.5-Turbo’s limitations in the unaligned case in Figure 4, Section F.6 provides a more general exploration of how well our constitutions can transfer between models. To this end, we take the best performing constitution of the unaligned case on the training set and test how well models from Anthropic’s Claude family are able to use these constitutions on the test set. The results indicate that this constitution transfers well to the Claude models — better than to GPT-3.5-Turbo, although transfer still incurs some loss. This is a promising result as it indicates that our constitutions are not excessively overfitting to the generation model, which indicates that they may capture more general concepts that are also interpretable to humans.
Scaling. Finally, although we consider sample-efficiency to be a benefit of our approach, we further evaluate whether these results also hold with larger scale datasets. We repeat the experiment on the AlpacaEval unaligned dataset with the full 648 preference pairs in the original dataset, using 324 samples each for training and testing. We observe that the overall results are very similar: the constitutional annotator with 61% agreement (66% prev.) still notably outperforms the default annotator with 34% (27% prev.). The full results are included in Section F.7.
@@ 4.3 Individual preferences: Chatbot Arena data
We evaluate ICAI’s ability to generate personal constitutions using the Chatbot Arena Conversations dataset (Zheng et al., 2023), which consists of 33k human-annotated preferences with user-generated prompts, offering richer insights into individual preferences compared to AlpacaEval. We select two users exhibiting different preferences from the Chatbot Arena dataset (with 9 and 12 preference annotations respectively) and generate a separate constitution with 3 principles for each. Due to limited samples, there is no training-test split. We provide details on the experimental setup and results in Section F.1 and summarize the results here. As may be expected, we find that the personal constitutions generated for each user are able to improve the annotator’s performance on the user’s annotations, but do not transfer well to the other user’s annotations. This shows that our constitutions successfully capture individual differences, illustrating the potential to generate personal constitutions. Note that how well a personal constitution transfers from one user to another depends on how similar their preferences are, with the contrast being the most pronounced between users with opposing preferences. Results will therefore vary for any given pair of users.
4.4 Demographic group preferences: PRISM data
Figure 5: Case-study: Constitutions for demographic groups on PRISM data. We consider two groups reported by Kirk et al. (2024) to have preferences differing from average: participants born in one geographical region rank Mistral-7b higher in this dataset (Group A), and those born in another region rank Llama-2-7b lower than average (Group B). We generate constitutions for both groups to explore these preferences. For each group, the annotator using the group’s data performs best. Constitutions (see Section H.4) suggest that Group A prefers Mistral-7b due to it’s conciseness, while Group B’s constitutions have recurring rules related to providing more detailed descriptions. Plots show mean and standard deviation (6 seeds) using GPT-4o.
Finally, we test ICAI’s ability to help interpret demographic group preferences on the PRISM dataset (Kirk et al., 2024), consisting of annotations on 8,011 multi-turn conversations with 21 LLMs across a range of value-based and controversial topics from a diverse set of 1,396 annotators.
Experimental setup. We consider two annotator subgroups described by Kirk et al. (2024) to have differing preferences with respect to certain models relative to the average annotators. Group A, annotators born in the same geographical region, are reported to prefer the outputs of Mistral-7b relative to the overall rank. Similarly, Group B, annotators born in a different region, are reported to dislike Llama-2-7b disproportionately. These observations raise the question: why do these subgroups prefer or dislike those specific models? With no concrete explanation in the original paper, ICAI offers a method to generate and test possible explanations. We select the subset of PRISM interactions where Group A prefers Mistral-7b over another model (30 pairs [^3] For each relevant PRISM datapoint, we pick one of three available rejected models at random.), and similarly, the subset where Group B rejects Llama-2-7b for any other model (80 pairs). Note that, as for Chatbot Arena, the few samples do not allow for a train/test split, which we consider acceptable for the purposes of data explanation. We then run ICAI on each subset separately to create a constitution for each group, testing each on both datasets. We use the same prompting setup as the Chatbot Arena experiments.
Results. Figure 5 shows that our constitutional annotators exceed default annotator performance in reconstructing the datasets they aim to compress but (as expected) do not transfer well to the other group’s annotations, in line with the observation by Kirk et al. (2024) that each group’s preference differs from average preferences. Indeed, the generated constitutions (see Section H.4) allow us to also ask how the preferences differ: Group A appears to strongly prefer more concise responses, whereas Group B has more diverse constitutions that often ask for more detailed descriptions.
4.5 Application: Bias detection
We showcase ICAI’s application in bias detection, following three steps: (1) Generate and test 400 candidate principles on 1,000 preference pairs (500 from PRISM, 500 from Chatbot Arena [^4] This experiment uses the Kaggle data (see Appendix A), differing from the data used in other experiments.) to capture diverse biases (using ICAI algorithm steps 1 to 4). (2) Manually select principles indicating potential biases, focusing on those with high accuracy or limited applicability. (3) Evaluate these principles on 13k preferences ($7{,}490$ from PRISM, $5{,}115$ from Chatbot Arena, and $648$ from AlpacaEval), using step 4 of the ICAI algorithm. All steps use GPT-4o-mini for cost efficiency.
Table 1: Evaluation of possibly biased principles on three datasets. Results on AlpacaEval ($648$ preferences), Chatbot Arena ($5{,}115$), and PRISM ($7{,}490$), showing relevance (fraction of data points where the principle applies) and accuracy (correctly reconstructed relevant data points). Grey values indicate relevance for fewer than $50$ preferences. See Table 2 in Section F.2 for extended results.
Principle AlpacaEv. ChatbotAr. PRISM
Select the response that… Acc Rel Acc Rel Acc Rel
is overly lengthy and lacks brevity 52.2 80.4 57.1 97.1 57.7 96.9
provides a numbered list format 73.4 12.2 62.3 45.5 71.6 17.7
presents a definitive stance without nuance 58.6 10.8 57.1 11.4 49.1 31.2
is overly general and vague 30.4 15.7 26.8 9.6 26.1 23.1
emphasizes neutrality over providing information 50.0 2.2 40.8 10.2 58.0 46.2
Results, shown in Table 1, reveal biases regarding verbosity, style, and assertiveness. Verbosity bias, where longer responses are preferred, is consistently observed across datasets, with Chatbot Arena and PRISM strongly favouring overly lengthy responses (notably, 57.1% and 57.7% accuracy, respectively). Style biases, such as a preference for numbered lists, are especially prominent in AlpacaEval (73% accuracy) and PRISM (72%), although their relevance is more limited compared to verbosity-related principles. Additionally, biases around ambiguity and vagueness vary by dataset; for example, PRISM annotations often favour neutral responses, while Chatbot Arena appears to actively select against response emphasizing neutrality. These results emphasize the dataset-specific nature of biases and demonstrate our framework’s sensitivity to such patterns. More biases, discussion and mitigation strategies can be found in Section F.2. Further, we provide an additional application example of annotation scaling on helpful/harmless data in Section F.3.
4.6 Ablation studies
We conduct ablation studies to assess the contribution of each step in our pipeline across four scenarios: synthetic orthogonal, synthetic aligned, synthetic unaligned, and AlpacaEval unaligned, using GPT-3.5-Turbo for the first three and GPT-4 for the last. Numerical results, as well as a detailed discussion of the ablations, the experimental setup, and the results, are provided in Section F.4. Key findings are summarized below:
Simplified principle generation (Step 1).
Generating only a single principle with a neutral prompt slightly reduces performance, indicating the importance of diverse principles. The performance drop is particularly pronounced on the synthetic aligned dataset, which aligns with our observation that GPT-3.5-Turbo struggles to generate both positive and negative principles from a single prompt.
No deduplication (Steps 2, 3, and 5).
Ablating deduplication produces mixed results: performance decreases on the synthetic aligned and synthetic unaligned datasets but improves on the synthetic orthogonal and AlpacaEval unaligned datasets. We hypothesize that repetition reinforces principles, especially when the model strongly opposes certain principles, as seen in AlpacaEval. Conversely, deduplication proves more effective when principles are less opposed to model biases, as observed in the synthetic orthogonal scenario. We conclude that deduplication is likely generally beneficial but that repetition may help overcome strong model biases in specific cases. Further details can be found in Section F.4.1.
No filtering and testing (Steps 4 and 5).
Removing the filtering and testing steps has the most drastic impact, with performance dropping across all datasets. In particular, the annotators fail to outperform the random baseline in the unaligned experiments.
@@ 5 Related work
Our work focuses on deriving interpretable principles from human feedback data and using AI annotators to evaluate those principles. We build on work related to learning from human feedback, biases in feedback, interpretable preference models, and AI annotators.
@@
Learning from human feedback. Fine-tuning LLMs with human feedback has significantly contributed to the success of modern LLMs (Ouyang et al., 2022; Stiennon et al., 2020). Typically, feedback is collected through pairwise comparisons of model outputs, training a reward model for fine-tuning, e.g. using reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Stiennon et al., 2020; Kaufmann et al., 2024) or direct preference optimization (DPO) (Rafailov et al., 2023). Interpreting preference data is challenging since it generally lacks annotations of the underlying reasons and the reward model is often a black-box neural network, making it hard to interpret. Our work aims to generate interpretable principles explaining the feedback data.
@@
Biases in human feedback. Identifying biases in human feedback data is crucial, as unintended biases are common. For example, Hosking et al. (2024) note a preference for assertiveness over truthfulness, while Wu & Aji (2023) highlight a bias towards grammatical correctness over factual accuracy. We discuss further prior work on annotator bias in Section B.2, also considering AI annotator bias. While these studies provide valuable insights, most methods for detecting biases rely on specialized feedback data collection, making them challenging to apply to pre-existing data. Our work generates interpretable principles from existing preference data, which can be inspected to detect biases and provide insights into the underlying preferences.
LLM-based description of dataset differences. Zhong et al. (2022; 2023) use LLMs to describe the difference between two text distributions in natural language, Dunlap et al. (2024b) a similar methodology for image distributions. Our work uses a related approach but adapted to the specific requirements and data structure of the pairwise preference domain.
@@
Interpretable preference models. There has been a growing interest in creating interpretable preference models, aiding in understanding behaviour of AI systems. Go et al. (2024) create a compositional preference model based on 13 fixed features, similar to our constitutional principles. While they do not generate the constitution from data, they do create a regression model to weigh them, which would be a promising extension to our approach. Petridis et al. (2024) propose a feedback-based constitution generation method relying on interactive tools, whereas our approach can be directly applied to standard preference datasets.
@@
AI annotators. Due to the cost and time required for human feedback collection, AI annotators, or LLM-as-a-judge, have been proposed as a scalable alternative. Constitutional AI (Bai et al., 2022b) uses LLMs with a set of principles for feedback. Human preference-based fine-tuning enables LLMs to generalize from very general principles, such as “do what’s best for humanity” (Kundu et al., 2023), or even give feedback well-aligned with human preferences without any constitution (Zheng et al., 2023). Our experiments show similar trends, where default LLM annotators align well with dataset annotations, even without a constitution. AI annotators can also exhibit biases, further discussed in Appendix B. AlpacaEval (Li et al., 2024c) offers a set of well-validated AI annotators.
@@
Rule-based preference learning. Rule learning, aiming to develop descriptive or predictive rules, has previously been applied to preference learning (de Sá et al., 2011). A common technique for rule learning is to first generate a set of candidate rules and then measure each rule’s support in a dataset, i.e. the fraction of data points that satisfy the rule (Fürnkranz et al., 2012; de Sá et al., 2011). Our algorithm follows this approach but, in contrast to more traditional rule learning, generates rules as natural language sentences. These rules, though more ambiguous and requiring AI annotators for interpretation, are expressive, interpretable, and easy for non-experts to edit [^5] This can also be seen as a method for automatic prompt generation, as discussed in Appendix B.. Liu et al. (2023) follow a similar generate-and-test approach to derive text quality criteria. They require absolute scores on a fixed set of aspects, however, while our method leverages pairwise comparisons covering many aspects simultaneously, as is commonly the case for publicly available preference datasets.
Concurrent work. Further, we would like to highlight concurrent works by Kostolansky (2024), Kostolansky & Manyika (2024), Shankar et al. (2024b) and Dunlap et al. (2024a) exploring ideas highly related to our work, perhaps highlighting the timeliness of this line of research. Whilst related, our work differs in terms of the precise choice of approach taken as well as the comprehensiveness of our experiments. We provide a detailed comparison in Section B.3.
6 Limitations